On average, a hectare produces 50 tons of bananas. For sake of simplicity, lets assume that they are only 100 hectares of bananas. Total production is then 5000 tons. Assume that disease will destroy 10% of the initial area in a year. It means that in year 0 production will be 5000 tons. In year 1 , production is 4500 tons. In year 2, production is 3500 tons .Assume the initial price is $1. Assume demand elasticity is -0.5 and supply elasticity is 1. Calculate KA for each year. Note that K will be calculated based on the change of total quantity produced instead of the change in yield. You will find a negative KA which is the vertical shift generated by the disease. How does this look graphically ? Draw the graph. 2. Calculate change in consumer and producer surplus change. Note that it is not positive. Assume that the discount rate is zero.ย 


1
Expert's answer
2021-11-10T10:07:17-0500

a.both years

"\frac{1}{0.5}=-0.2"


b.



Producer surplus=5000+3500=1500

consumer surplus=4500-3500=1000